# Unmasking Backdoors: An Explainable Defense via Gradient-Attention Anomaly Scoring for Pre-trained Language Models

### v1

@@ 1 Introduction
Most existing defense strategies center around identifying and eliminating malicious samples either during training stage or at inference. Training-time defenses Li et al. (2021b); Chen & Dai (2021) typically demand exhaustive monitoring of the entire dataset to identify and discard poisoned samples. Alternatively, some methods necessitate splitting the dataset into several partitions and training multiple models concurrently, a strategy that significantly escalates computational costs Pei et al. (2023). This condition is especially hard to fulfill in typical pre-train–fine-tune workflows, as pre-training is usually handled by third parties. Some defense methods Shen et al. (2022) aim to detect backdoored models and deploy trigger inversion strategies to recover and neutralize the injected triggers. These methods are computationally expensive for large-scale applications, as they require an extensive search over the token space. Inference-time defenses Qi et al. (2021a) attempt to filter poisoned samples during prediction via auxiliary detection workflows. However, identifying such inputs is challenging due to the unknown nature of attacker-injected triggers Yang et al. (2021c); Qi et al. (2021b). Moreover, most existing defenses offer little or no insight in terms of explainability.
@@ 2 Related Work
Defense. Existing backdoor defense strategies primarily focus on detecting and removing poisoned samples either during training or inference, or on purifying the compromised model to eliminate embedded backdoors. Training-time defenses often rely on trigger pattern analysis Chen & Dai (2021) or clustering-based methods Cui et al. (2022). Some approaches partition the dataset and train multiple models in parallel to isolate poisoned behavior Pei et al. (2023). Inference-time defenses, on the other hand, typically employ techniques such as perplexity-based filtering Qi et al. (2020) or rare word perturbation Yang et al. (2021b). Model purification-based defenses aim to eliminate backdoors from compromised models using various strategies. Some approaches rely on the availability of a clean model and merge its weights with those of the backdoored model Zhang et al. (2022; 2023). Others, such as MEFT Liu et al. (2023), introduce maximum entropy training to suppress malicious behavior before fine-tuning on clean data. Another approach, PURE Zhao et al. (2024b), focuses on the attention variance of the [CLS] token. It detects and prunes heads with low [CLS] attention variance on clean inputs, normalizes attention distributions, and subsequently fine-tunes the model to mitigate backdoor effects. However, these strategies are often computationally expensive, requiring retraining, fine-tuning, or exhaustive token space exploration, and typically lack interpretability. In contrast, our method is an inference-time defense that leverages token-level attributes-attention and gradient signals to construct anomaly scores. This enables both backdoor mitigation and explainable insights into the location and behavior of trigger tokens.
@@ 3.2 Threat Model
Defender’s Goal and Capabilities. Given a backdoored model $\tilde{M}$, the defender aims to reduce the attack success rate while preserving the model’s classification accuracy on clean data. The defender has access to a small set of clean validation samples, denoted as $\mathcal{D}_{clean}^{val}\sim\mathcal{D}$. However, the defender possesses no prior knowledge about the backdoor triggers or the target label(s) chosen by the attacker.
@@ 5.2 Results and Analysis
Figure 3: Robustness evaluation under varying anomaly thresholds. Top: LWS attack on SST-2, Bottom: BADNET attack on IMDb.
To assess the robustness of our defense framework, we evaluate its sensitivity to different threshold percentiles used for anomaly detection. Lowering the threshold results in more samples being flagged as suspicious and undergoing token corruption, which in turn increases computation time. This typically improves ASR, but may risk degrading CACC, as some clean samples might be unnecessarily perturbed. However, as shown in Fig. 3, our findings indicate that CACC remains largely stable across a wide range of thresholds for BERT and DistilBERT, indicating that the noise injection module minimally affects clean samples. Notably, for ALBERT on the SST-2 dataset under the LWS attack, we observe that as the threshold is lowered (from 95% to 45%), ASR steadily decreases, as expected, while CACC surprisingly increases. This suggests that the corruption applied by the Trigger Neutralizer suppresses backdoor activation and also pushes poisoned predictions toward the correct class, without adversely impacting clean samples, further validating the effectiveness and precision of our proposed defense.
@@ A.1.1 Evaluation Under Domain Shift (Task-Agnostic Setting)
A.1.2 Token Attribution Patterns in Backdoored Inputs
@@
To better understand the influence of individual tokens under backdoor attacks, we visualize the token attribution scores computed using Eqn. 7 on the SST-2 dataset. Figure 5(a) shows the attribution scores for a representative poisoned input: ”it’s hampered by a lifetime-channel kind of plot and a lead actress who is tq out of her depth.” Here, the trigger token tq clearly stands out with a significantly higher score compared to benign tokens. Moreover, Figure 5(b) presents the average attribution scores aggregated over all poisoned samples. This global view reveals that certain tokens (e.g., mn, mb, tq) consistently receive disproportionately high scores, confirming their role as backdoor triggers. These findings highlight the interpretability of our attribution-based scoring mechanism in localizing suspicious trigger tokens and explaining model behavior under backdoor attacks.
A.1.3 Evaluation on RoBERTa and DeBERTa Models
@@ A.5 Implementation Details of Backdoor Defense
We summarize the implementation details of our defense approach. As described in the main paper, we evaluate our method on two representative NLP tasks: sentiment analysis and topic classification, using the SST-2, IMDb, and AG’s News datasets. For each dataset, we reserve 20% of the original training data to construct the clean validation set $\mathcal{D}_{\text{clean}}^{\text{val}}$, which is used to determine the anomaly score threshold ($\tau$). The threshold $\tau$ is set to the 95th percentile of the anomaly scores computed on a clean validation set for BERT, RoBERTa, DistilBERT, and DeBERTa. For ALBERT, we adopt a lower threshold at the 65th percentile, as we observe that while backdoor trigger tokens continue to dominate in token-level attribution scores, the overall sentence-level anomaly scores tend to be lower. All experiments are implemented in the PyTorch framework using the Hugging Face Transformers library. We perform the defense evaluations on a single NVIDIA A100 GPU.

### v2

@@ 1 Introduction
Most existing defense strategies center around identifying and eliminating malicious samples either during training stage or at inference. Training-time defenses Li et al. (2021b); Chen and Dai (2021) typically demand exhaustive monitoring of the entire dataset to identify and discard poisoned samples. Alternatively, some methods necessitate splitting the dataset into several partitions and training multiple models concurrently, a strategy that significantly escalates computational costs Pei et al. (2023). This condition is especially hard to fulfill in typical pre-train–fine-tune workflows, as pre-training is usually handled by third parties. Some defense methods Shen et al. (2022) aim to detect backdoored models and deploy trigger inversion strategies to recover and neutralize the injected triggers. These methods are computationally expensive for large-scale applications, as they require an extensive search over the token space. Inference-time defenses Qi et al. (2021a) attempt to filter poisoned samples during prediction via auxiliary detection workflows. However, identifying such inputs is challenging due to the unknown nature of attacker-injected triggers Yang et al. (2021c); Qi et al. (2021b). Moreover, most existing defenses offer little or no insight in terms of explainability.
@@ 2 Related Work
Defense. Existing backdoor defense strategies primarily focus on detecting and removing poisoned samples either during training or inference, or on purifying the compromised model to eliminate embedded backdoors. Training-time defenses often rely on trigger pattern analysis Chen and Dai (2021) or clustering-based methods Cui et al. (2022). Some approaches partition the dataset and train multiple models in parallel to isolate poisoned behavior Pei et al. (2023). Inference-time defenses, on the other hand, typically employ techniques such as perplexity-based filtering Qi et al. (2020) or rare word perturbation Yang et al. (2021b). Model purification-based defenses aim to eliminate backdoors from compromised models using various strategies. Some approaches rely on the availability of a clean model and merge its weights with those of the backdoored model Zhang et al. (2022; 2023). Others, such as MEFT Liu et al. (2023), introduce maximum entropy training to suppress malicious behavior before fine-tuning on clean data. Another approach, PURE Zhao et al. (2024b), focuses on the attention variance of the [CLS] token. It detects and prunes heads with low [CLS] attention variance on clean inputs, normalizes attention distributions, and subsequently fine-tunes the model to mitigate backdoor effects. However, these strategies are often computationally expensive, requiring retraining, fine-tuning, or exhaustive token space exploration, and typically lack interpretability. In contrast, our method is an inference-time defense that leverages token-level attributes-attention and gradient signals to construct anomaly scores. This enables both backdoor mitigation and explainable insights into the location and behavior of trigger tokens.
@@ 3.2 Threat Model
Defender’s Goal and Capabilities. Given a backdoored model $\tilde{M}$, the defender aims to reduce attack success rate while preserving clean accuracy. The defender has access to a small set of clean validation samples, denoted as $\mathcal{D}_{clean}^{val}\sim\mathcal{D}$. However, the defender possesses no prior knowledge about the backdoor triggers or the target label(s) chosen by the attacker. Following prior work Liu et al. (2023); Zhao et al. (2024a), we assume a white-box pre-train–fine-tune setting with access to model parameters, gradients, and attention weights, mirroring the attacker’s visibility.
@@ 5.2 Results and Analysis
Figure 3: Sensitivity under varying anomaly thresholds. Top: LWS attack on SST-2, Bottom: BADNET attack on IMDb.
We evaluate its sensitivity to different threshold percentiles used for anomaly detection. Lowering the threshold results in more samples being flagged as suspicious and undergoing token corruption, which in turn increases computation time. This typically improves ASR, but may risk degrading CACC, as some clean samples might be unnecessarily perturbed. However, as shown in Fig. 3, our findings indicate that CACC remains largely stable across a wide range of thresholds for BERT and DistilBERT, indicating that the noise injection module minimally affects clean samples. Notably, for ALBERT on the SST-2 dataset under the LWS attack, we observe that as the threshold is lowered (from 95% to 45%), ASR steadily decreases, as expected, while CACC surprisingly increases. This suggests that the corruption applied by the Trigger Neutralizer suppresses backdoor activation and also pushes poisoned predictions toward the correct class, without adversely impacting clean samples, further validating the effectiveness and precision of our proposed defense.
@@ A.1.1 Evaluation Under Domain Shift (Task-Agnostic Setting)
A.1.2 Evaluation Under Clean-Label Backdoor Attacks
To assess the generalizability of X-GRAAD, we also evaluate it under clean-label backdoor attacks, where poisoned samples preserve their original ground-truth labels. Following the clean-label protocol of Cui et al. (2022), we use the rare-word trigger (cf) and apply a 20% poisoning ratio on SST-2 to backdoor a BERT model. This setup successfully implants a clean-label backdoor, yielding ASR = 1.0 and CACC = 0.93, while accuracy on poisoned samples drops to 0.476. After applying X-GRAAD, the ASR falls drastically from 1.0 to 0.012, and accuracy improves from 0.476 to 0.910 (CACC = 0.915). These results indicate that the attention–gradient anomaly signal isolates the rare-word trigger, demonstrating that X-GRAAD generalizes effectively to clean-label, trigger-based backdoors without any modification to the defense pipeline.
A.1.3 Token Attribution Patterns in Backdoored Inputs
@@
To better understand the influence of individual tokens under backdoor attacks, we visualize the token attribution scores computed using Eqn. 7 on the SST-2 dataset. Fig.  5(a) shows the attribution scores for a representative poisoned input: ”it’s hampered by a lifetime-channel kind of plot and a lead actress who is tq out of her depth.” Here, the trigger token tq clearly stands out with a significantly higher score compared to benign tokens. Moreover, Fig.  5(b) presents the average attribution scores aggregated over all poisoned samples. This global view reveals that certain tokens (e.g., mn, mb, tq) consistently receive disproportionately high scores, confirming their role as backdoor triggers. These findings highlight the interpretability of our attribution-based scoring mechanism in localizing suspicious trigger tokens and explaining model behavior under backdoor attacks.
A.1.4 Evaluation on RoBERTa and DeBERTa Models
@@
A.1.5 Robustness Analysis
(a) ASR and CACC with/without defense for three triggers.
(b) Token attribution for ”the silver quill”.
(c) Token attribution for ”a crimson echo”.
(d) Token attribution for ”james bond”.
Figure 6: (a) Effect of X-GRAAD on attack ASR and CACC for multi-token distributed triggers; (b–d) token-level attribution on single poisoned samples.
Table 4: Performance of various defenses on multi-token backdoor triggers under the BadNets attack on BERT (SST-2).
Trigger Dataset SST-2
Method ONION RAP FT MEFT PURE X-GRAAD
”the silver quill” BadNets CACC 0.853 0.912 0.931 0.925 0.921 0.909
ASR 0.317 1.000 0.040 0.039 0.025 0.012
”a crimson echo” BadNets CACC 0.861 0.919 0.931 0.926 0.920 0.906
ASR 0.204 1.000 0.995 0.985 0.850 0.058
”james bond” BadNets CACC 0.861 0.913 0.937 0.935 0.918 0.914
ASR 0.135 0.81 0.198 0.358 0.039 0.022
We conduct an additional robustness analysis on BERT for the SST-2 dataset under the BadNets attack to examine how X-GRAAD behaves when the backdoor trigger spans multiple tokens. For this, we choose short, rare, yet semantically meaningful phrases that satisfy the stealthy rare-word properties typical of practical backdoor triggers (e.g., ”the silver quill”, ”a crimson echo”, ”james bond”). The results are presented in Table 4. Although these triggers span multiple tokens, we observe that the poisoned model consistently collapses the backdoor dependency onto a single high-saliency pivot token (Fig. 6(b)6(d)), which X-GRAAD reliably identifies using the max-based anomaly score.
As shown in Fig. 6(a), X-GRAAD substantially reduces ASR while preserving clean accuracy: for ”the silver quill”, ASR decreases from 1.00 to 0.012 (CACC 0.918 $\rightarrow$ 0.909); for ”a crimson echo”, from 1.00 to 0.058 (0.920 $\rightarrow$ 0.906); and for ”james bond”, from 0.83 to 0.022 (0.924 $\rightarrow$ 0.914). These results confirm that multi-token triggers remain detectable and neutralizable under our framework. Token-level attribution further supports the pivot phenomenon: crimson (7.71), bond (33.01), and qui/ll (22.73 / 10.29) dominate the anomaly scores despite appearing within multi-token phrases (Fig. 6(b)6(d)).
@@ A.5 Implementation Details of Backdoor Defense
We summarize the implementation details of our defense approach. As described in the main paper, we evaluate our method on two representative NLP tasks: sentiment analysis and topic classification, using the SST-2, IMDb, and AG’s News datasets. For each dataset, we reserve 20% of the original training data to construct the clean validation set $\mathcal{D}_{\text{clean}}^{\text{val}}$, which is used to determine the anomaly score threshold ($\tau$). The threshold $\tau$ is set to the 95th percentile of the anomaly scores computed on a clean validation set for BERT, RoBERTa, DistilBERT, and DeBERTa. This follows the standard mean + 2 standard-deviation principle (95th) commonly used in anomaly detection. For ALBERT, whose shared-layer architecture; a single attention block reused across all layers; produces a noticeably compressed attribution distribution, since it lacks the diversity of signals present in multi-layer encoders such as BERT or DistilBERT. As a result, ALBERT’s anomaly scores are systematically lower and less dispersed, and a slightly lower threshold is required to separate clean and poisoned samples. We therefore use the 65th percentile, which closely aligns with the standard mean + 1 standard-deviation (68th percentile) rule. All experiments are implemented in the PyTorch framework using the Hugging Face Transformers library. We perform the defense evaluations on a single NVIDIA A100 GPU.